the length of a rectangle is 9 in.greater than the width. The area is 36 in.2

Find the dimensions.

Answer by checkley77(12844)

L=W+9
LW=36
(W+90)W=36
W^2+9W-36=0
(W+12)(W-3)=0
W-3=0
W=3 ANS. FOR THE WIDTH.
L=3+9
L=12 ANS. FOR THE LENGTH.
PROOF:
3*12=36
36=36